# A Survey of Secure Semantic Communications

### v1

@@ 3.1 Life Cycle of SemCom
Figure 4: Life cycle of SemComs.
@@ 3.3.1 Knowledge Base Poisoning Attacks
Malicious clients can generate dirty samples to train the global SemCom model to produce falsified model parameters and send them to the server [85]. The most commonly used data poisoning strategy is label-flipping [88], where the label of training samples is modified to another category, resulting in taking control of the client’s local models and ultimately manipulating the global model [85]. The objective of the data poisoning attacks is generally to degrade the accuracy of learning tasks, increase the convergence time of the global model, and the probability of erroneous learning results [89].
@@ 3.4 Security and Privacy Issues in Model Transfer Stage
Constrained by the network capacity, the AI model needs to be broken up into smaller fragments for transport and then reassembled when entering the receiving node, a process usually known as fragmentation [99]. Inspired by the works to deal with AI-related tasks in real-time with bandwidth, computational capability, memory and delay limit via splitting the model [105, 106, 107, 108], model partitioning is adopted in SemCom systems. In detail, the model is partitioned as a sequential connection of several layers and then encapsulated into packets for transmission throughout the network.
3.4.1 Model Slice Attacks
During the transmission of model slices [61], if the isolation measures between them are not sufficiently rigorous, attackers may capitalize on this vulnerability to steal sensitive information contained within the slices, including model parameters and structures, ultimately launching attacks on the model. Moreover, uneven resource allocation can potentially spark DoS attacks, where attackers block slice transmission channels by consuming substantial resources, thereby preventing model slices from reaching the receiving node as intended and disrupting the overall system operation [109]. Furthermore, virtual network functions are susceptible to risks of unauthorized access and tampering [110]. If attackers successfully penetrate the network’s security perimeter and illegally access or tamper with the slices, it will severely undermine the security and confidentiality of the slice environment, resulting in the failure of the model to be accurately reassembled or executed at the receiving node.
@@ 3.5.1 Semantic Adversarial Attacks
Adversarial attacks mislead learning models by subtly altering input data [112]. These attacks are equally pertinent to SemCom systems, which rely on DL models for the extraction, encoding, and transmission of semantic information. Semantic adversarial attacks are covert and efficient [24]. Attackers can intentionally introduce imperceptible perturbations into the input data, causing the model to make erroneous predictions or exhibit unintended behaviors. These perturbations are barely noticeable to human senses but can exert significant impacts on DL models. Additionally, during the process of SemCom, wireless channels may be susceptible to various interferences and noises, creating vulnerabilities for attackers [113]. By introducing meticulously crafted perturbations into the channels, attackers can disrupt the transmission and reception of semantic information.
@@ 3.5.2 Semantic Eavesdropping Attacks
Wireless channels, which serve as the conduit for signals traveling from the sender to the receiver, are characterized by their openness. This implies that during transmission, signals are not solely received by legitimate receivers but can also potentially be intercepted by nearby devices operating within the same frequency band. This inherent openness makes wireless channels vulnerable to a multitude of noises, interferences, and other unpredictable time-varying disruptions, including those arising from user mobility [114]. Eavesdroppers can capitalize on this openness by employing specialized receiving equipment to capture and analyze signals in transit. Such equipment may encompass high-sensitivity antennas, signal amplifiers, demodulators, and various other tools dedicated to receiving, amplifying, and dissecting wireless signals. In certain instances, eavesdroppers may even deploy sophisticated signal processing techniques, such as signal separation and interference suppression, to boost the success rate and precision of their eavesdropping endeavors. If eavesdroppers obtain access to the semantic decoder, they could potentially decode some or all of the private or sensitive semantic information, even when their channel conditions are substantially inferior to those of the legitimate receivers [36].
@@ 3.5.3 Semantic Inference Attacks
The semantic inference attacker acquires sensitive information pertaining to the training data through queries to the DL model. Despite the fact that the model does not directly disclose the training data, by scrutinizing its outputs, the attacker can deduce attributes related to the training data, encompassing sample distribution, class information, or even individual training samples. Inference attacks encompass a range of tactics, including membership inference attacks, attribute inference attacks, and quantity inference attacks, among others.
@@ 3.5.4 Semantic Jamming Attacks
Table 3: The detailed relationship between security and privacy threats and defensive techniques.
Methods Descriptions Attacks Positions of Attacks Phases of Attacks Protected Properties
Data cleaning Data cleaning is used to deal with dirty data that may introduce noise, missing values or inconsistencies Knowledge base poisoning attacks Knowledge base Model training Integrity and Availability
Robust learning Robust learning is used to improve the resistance of DL models to noise or input disturbance, thus ensuring the stability and reliability Knowledge base poisoning attacks Knowledge base Model training Integrity and Availability
Defensive methods of backdoor attacks Both on the data and model level Backdoor attacks Knowledge base and source of information Model training and semantic information transmission Integrity and Availability
Adversarial training Adversarial training is realized by adding adversarial attacks to the training samples Semantic adversarial attacks and semantic inference attacks Source of information and physical channel Semantic information transmission Integrity and Availability
Differential privacy Differential privacy interferes with the original data by introducing noise into the data, making it impossible for attackers to accurately restore any individual information from the interfered data Gradient leakage and semantic inference attacks Model transmission Model training and semantic information transmission Confidentiality
@@
Cryptography technology Homomorphic encryption, secure multi-party computation, trusted execution environment, and secure aggregation Vulnerability of servers, model slice attacks, semantic eavesdropping attacks, and semantic jamming attacks Model transmission, physical channel, and receiver Model training, model transfer, and semantic information transmission Confidentiality, Integrity, and Availability
Block-chain technology Blockchain combine a series of data blocks into a chain data structure in a sequential way, and ensures that these data blocks cannot be tampered with and forged Model slice attacks, semantic adversarial attacks, semantic eavesdropping attacks, semantic inference attacks, and semantic jamming attacks Model transmission, physical channel, and receiver Model transfer and semantic information transmission Confidentiality and Integrity
Model compression Model pruning, parameter quantization, low-rank decomposition, and knowledge distillation Attacks against communication bottlenecks, semantic inference attacks Model transmission Model training, model transfer, and semantic information transmission Integrity and Availability
@@
Physical-layer security Beamforming, artificial noise, relay cooperation, intelligent reflecting surface, physical-layer key generation, and physical-layer authentication Semantic adversarial attacks, semantic eavesdropping attacks, semantic inference attacks, and semantic jamming attacks Physical channel and receiver Semantic information transmission Confidentiality, Integrity, and Availability
@@ 4 Techniques to Mitigate Security and Privacy Threats in SemCom
In response to the security and privacy threats detailed in Sec. 3, this section provides defensive techniques, including data cleaning, robust learning, defensive methods of backdoor attacks, adversarial training, differential privacy, cryptography technology, block-chain technology, model compression, and physical-layer security. The detailed relationship between them is shown in Tab. 3.
@@ 4.1 Data Cleaning
Dirty data can introduce noise, missing values, or inconsistencies, directly causing semantic misunderstandings and even communication failure [117]. Therefore, the effectiveness of data cleaning is particularly important. Fig. 5 illustrates the use of data cleaning in DL.
Figure 5: Data cleaning in DL [118].
The below, based on multiple academic studies, delves into various data cleaning methods for different data types and their applications in SemCom.
@@ 4.1.2 Diverse Cleaning of Image Data
Image data cleaning refers to the process of reducing noise, correcting artifacts, and enhancing data quality in images used for semantic communication. This includes techniques like noise removal through Gaussian blur and bilateral filtering, which preserve critical features while eliminating distortions. Data augmentation methods, such as rotations and scaling, further improve adaptability to diverse visual scenarios, and GANs can generate synthetic data to address training data shortages.
@@ 4.1.4 Comprehensive Optimization of Structured Data
Lesson 1
Data cleaning is essential for SemCom systems as dirty data can lead to semantic misunderstandings and communication failure, and its effectiveness is highlighted by various data cleaning methods for different data types in SemCom, including textual [119, 120, 121], image [122, 123, 120], audio [117, 119], and structured data [121, 119, 120].
@@ 4.2 Robust Learning
The design characteristics of FL, as shown in Fig. 6, make FL face numerous challenges in practical applications, such as noise interference, data heterogeneity, malicious clients, and low communication efficiency. These challenges not only hinder the improvement of model performance but also pose higher demands on system robustness and efficiency. This paper reviews recent advances in the robustness of FL and proposes potential directions for improvement.
Figure 6: Diagram of FL [124].
@@ 4.2.3 Mitigating the Impact of Noise and Malicious Clients
Figure 7: Malicious client attacks in FL [128].
@@ 4.2.4 Balancing Robustness and Communication Efficiency
Figure 8: Comparison of noise distributions under the expectation-based model and the worst-case model [130].
Lesson 2

### v2

@@ 3.1 Life Cycle of SemCom
Figure 4: Life cycle of SemComs, where a centralized model is first trained on the cloud and then transferred to edge servers. At the edge, the model and its slices are exchanged among servers to enable a robust co-training process that involves local training, model updates, aggregation into a refined global model and sharing it. This updated model is then delivered to the end node, and any further update requests trigger a new cycle of collaborative re-training, ensuring continuous optimization and efficient semantic information transmission throughout the network.
@@ 3.3.1 Knowledge Base Poisoning Attacks
Malicious clients can generate dirty samples to train the global SemCom model to produce falsified model parameters and send them to the server [85]. The most commonly used data poisoning strategy is label-flipping [88], where the label of training samples is modified to another category, resulting in taking control of the client’s local models and ultimately manipulating the global model [85]. The objective of the data poisoning attacks is generally to degrade the accuracy of learning tasks, increase the convergence time of the global model, and the probability of erroneous learning results [89]. For example, Peng et al. [90] present the targeted clean-label poisoning attack (TCLA) in multiuser SemCom, where an adversary meticulously selects and alters a subset of training data. This alteration introduces subtle perturbations that cause the SemCom model to misclassify specific instances during inference. Unlike conventional poisoning attacks that involve label flipping or the introduction of noisy data, TCLA achieves its objective without modifying the actual ground truth labels. Instead, it inserts meticulously crafted samples into the training dataset that resemble the target instances the attacker wishes to misclassify. By carefully selecting these instances, the attacker influences the model’s decision boundaries for specific classes or instances, potentially leading to security or privacy risks.
@@ 3.4 Security and Privacy Issues in Model Transfer Stage
Constrained by the network capacity, the AI model needs to be broken up into smaller slices for transport and then reassembled when entering the receiving node, a process usually known as fragmentation [100]. Inspired by the works to deal with AI-related tasks in real-time with bandwidth, computational capability, memory and delay limit via splitting the model [106, 107, 108, 109], model slices are adopted in SemCom systems. Slicing can divide the model into multiple smaller parts, thereby reducing the data volume per transmission and increasing transmission speed. During the model slice transfer stage, it becomes easier to detect which slice has failed or contains errors, and only those faulty slices, rather than the entire model, need to be retransmitted, thus saving time and bandwidth. Therefore, slicing allows for flexible adjustment and expansion of the model according to different application scenarios and requirements. We believe that different levels of security are needed to safeguard different model slices. For example, the attacker can directly identify the ground-truth labels based on the transmitted gradients on the last layer of the classification model trained with cross-entropy loss during the model update process [96]. In other words, the model slices should no longer be treated as non-differential bits when transmitted through the physical channel, and each model slice is with different vulnerability level. Therefore, it is reasonable that the model slice with high vulnerability, such as the last layer of the cross-entropy-trained model, needs extra protection when transmitted.
However, during the model slice transfer stage, it is susceptible to availability attacks and forgery attacks. Availability attacks can corrupt the parameters of the model slices, directly impacting their SemCom performance once deployed. Forgery attacks involve deploying model parameters pre-trained by the attacker at the terminal, thereby steering the user’s SemCom results towards the direction desired by the attacker.
3.4.1 Model Slice Availability Attacks
During the semantic model slice transfer stage, each slice may contain specific parts of the model, such as parameters, structures, or training data. These slices must be strictly isolated and protected during transmission to prevent unauthorized access. If the isolation measures are not sufficiently strict, attackers may have the opportunity to steal sensitive information from these slices. Once attackers obtain this sensitive information, they can reconstruct the model or use it to launch targeted attacks. Additionally, if the slices contain training data or derived information, the leakage of these data may violate personal privacy, especially when dealing with sensitive data such as medical records and financial information. Besides isolation measures, resource allocation is crucial in the transmission of semantic model slices. Uneven resource allocation may lead to unfair treatment of certain slices during transmission, causing a series of problems. For example, Denial of Service (DoS) attackers can block slice transmission channels by consuming large amounts of resources (such as bandwidth and computing resources). When attackers successfully block the channels, the model slices cannot reach the receiving nodes as expected, resulting in the inability to fully reconstruct or update the model, and ultimately disrupting the operation of the entire system. Furthermore, uneven resource allocation may also lead to a decline in the overall system performance, with increased latency and reduced throughput. This reduces the availability and reliability of the system, making it susceptible to instability even under normal loads.
3.4.2 Model Slice Forgery Attacks
Semantic model slices, by aggregating data, computing power, and domain knowledge, can be directly applied to commercial services, such as semantic IoT, possessing clear economic value. This makes semantic models, as core intellectual property, a common practice for trading and sharing. However, attackers may attempt to obtain unauthorized semantic model slices by forging seemingly legitimate model slice requests. For instance, model slice forgery attackers can exploit loose access control policies or logical vulnerabilities in semantic slice models to bypass normal security mechanisms and directly access and download semantic model slices. Additionally, during the semantic model slices transfer stage, strong identity binding mechanisms (such as digital certificates, hardware fingerprints, etc.) are crucial for ensuring slice security and preventing illegal access. However, if the semantic slice model management system lacks these strong identity binding mechanisms, attackers can more easily forge identity information, impersonate legitimate users, and initiate illegal requests.
@@ 3.5.1 Semantic Adversarial Attacks
Adversarial attacks mislead learning models by subtly altering input data [111]. These attacks are equally pertinent to SemCom systems, which rely on DL models for the extraction, encoding, and transmission of semantic information. Semantic adversarial attacks are covert and efficient [24]. Attackers can intentionally introduce imperceptible perturbations into the input data, causing the model to make erroneous predictions or exhibit unintended behaviors. These perturbations are barely noticeable to human senses but can exert significant impacts on DL models. Additionally, during the process of SemCom, wireless channels may be susceptible to various interferences and noises, creating vulnerabilities for attackers [112]. By introducing meticulously crafted perturbations into the channels, attackers can disrupt the transmission and reception of semantic information. For example, Sagduyu et al. [113] present the process of semantic adversarial attacks as follows. During the inference stage, semantic adversarial attacks aim to tweak the input data fed into the adversary’s model. This tweaking, often achieved through minor alterations or perturbations, renders the model unable to make accurate decisions for those particular samples. The effectiveness of these attacks is evaluated by examining the model’s accuracy on the altered test inputs. The more the accuracy drops, the more successful the attack is deemed. The perturbations are chosen by minimizing their impact while ensuring that they cause the victim model to make errors. Additionally, the power of these perturbations is kept within a certain threshold. Adversaries can launch both targeted and non-targeted attacks. Targeted attacks aim to mislead the semantic model’s outputs specifically for a set of non-target labels, redirecting them to target labels by minimizing the loss function related to those target labels. Conversely, non-targeted attacks strive to cause errors for samples from all labels by maximizing the loss function for all samples under attack.
@@ 3.5.2 Semantic Eavesdropping Attacks
Wireless channels, which serve as the conduit for signals traveling from the sender to the receiver, are characterized by their openness. This implies that during transmission, signals are not solely received by legitimate receivers but can also potentially be intercepted by nearby devices operating within the same frequency band. This inherent openness makes wireless channels vulnerable to a multitude of noises, interferences, and other unpredictable time-varying disruptions, including those arising from user mobility [114]. Semantic eavesdroppers can capitalize on this openness by employing specialized receiving equipment to capture and analyze signals in transit. Such equipment may encompass high-sensitivity antennas, signal amplifiers, demodulators, and various other tools dedicated to receiving, amplifying, and dissecting wireless signals. In certain instances, semantic eavesdroppers may even deploy sophisticated signal processing techniques, such as signal separation and interference suppression, to boost the success rate and precision of their eavesdropping endeavors. If eavesdroppers obtain access to the semantic decoder, they could potentially decode some or all of the private or sensitive semantic information, even when their channel conditions are substantially inferior to those of the legitimate receivers [36].
@@ 3.5.3 Semantic Inference Attacks
By analyzing intermediate information exposed during model training, semantic inference attackers can infer whether specific data belongs to the training set members and even further reconstruct sensitive semantic information. This attack method directly threatens the privacy security of each user participating in distributed semantic learning. Specifically, attackers can achieve privacy threats by directly exploiting shared information, poisoning the training process, and conducting collaborative attacks as follows.
  • 1.
    Direct exploitation of shared semantic information: By observing the gradient distribution of the global model, attackers can identify the contribution of specific samples to the gradients. If the gradient direction of a sample is highly correlated with the global update, it can be inferred that the sample is likely part of the training set. Furthermore, semantic embeddings in shared knowledge bases may implicitly contain the distribution characteristics of training data. Attackers can use reverse engineering to infer original data from embedding semantic representations, for example, by discovering frequently occurring data categories through clustering analysis.
  • 2.
    Poisoning the training process: Attackers can upload local data containing adversarial samples, causing the model to overfit on specific semantic categories. This overfitting phenomenon amplifies the possibility of semantic inference because the introduction of adversarial samples alters the model’s training trajectory. Additionally, during local training, attackers can inject malicious gradients, such as by scaling gradients or shifting their directions, to interfere with global model parameters. This interference can induce legitimate users to expose more sensitive semantic information during the inference phase. Moreover, attackers can modify the loss function during local training, such as by adding penalties for specific samples, causing the model to behave abnormally when processing these data. This abnormal behavior also aids attackers in subsequent semantic inference.
  • 3.
    Collaborative attacks: Multiple malicious participants can jointly launch semantic inference attacks, enhancing the accuracy of inference by cross-referencing gradient or parameter updates from different participants. This collaborative semantic inference attack method can further amplify the risk of privacy leakage because multiple attackers can share information and work together to improve the attack effectiveness.
@@ 3.5.4 Semantic Jamming Attacks
Table 4: List of Defensive Techniques for Secure SemCom.
Methods Sub-Class Attacks Effectiveness
Data cleaning Data cleaning for textual data, image data, audio data, and structured data Knowledge base poisoning attacks It simplifies the data set by removing redundant, repetitive and invalid data, thus reducing the data processing burden in the process of semantic model training. Additionally, it can identify and correct errors and abnormal values in data sets, thus improving the accuracy of data and reducing misunderstandings and ambiguities caused by data errors.
Robust learning Robust aggregation, data heterogeneity, the impact of noise and malicious clients, and balancing robustness and communication efficiency Knowledge base poisoning attacks Robust learning can enhance the semantic model’s anti-interference ability to abnormal data and noise, so that the SemCom system can still maintain stable and reliable performance in the face of all kinds of interference and noise, and improve the semantic model’s resistance ability in the face of malicious attacks or data tampering.
Defensive methods of backdoor attacks Data processing and model identification and construction Backdoor attacks It can be achieved from data and model perspectives.
Adversarial training Adversarial sample generation, adversarial regularization, adaptive disturbance constraint, and matching pluggable modules Semantic adversarial attacks and semantic inference attacks Adversarial training is realized by adding adversarial attacks to the training samples to defend against semantic adversarial attacks. In SemCom, semantic adversarial samples are generated by adding noise to the semantic information to train the model together with the semantic information in the original training set.
Differential privacy Parameter value protection, objective function perturbation, and data generation and pre-processing Gradient leakage and semantic inference attacks In SemCom, the single semantic symbol cannot influence the meaning of the whole sentence. Therefore, the risk of privacy breaches arising from the addition of a piece of semantic symbol to a dataset is kept so small that an attacker cannot obtain accurate information by observing the results of the computation.
@@
Cryptography technology Homomorphic encryption, secure multi-party computation, trusted execution environment, secure aggregation, and quantum cryptography Vulnerability of servers, model slice availability attacks, model slice forgery attacks, semantic eavesdropping attacks, and semantic jamming attacks Encryption algorithm can convert plaintext data into ciphertext, which enhances the confidentiality of semantic information. Integrity verification algorithms can ensure that the semantic data has not been tampered with during transmission and confirm the authenticity and integrity of the semantic data. Cryptography technology can also be used for identity authentication to ensure the true identity of both parties.
Block-chain technology Data security, prevent node leakage, and model framework design Model slice availability attacks, model slice forgery attacks, semantic adversarial attacks, semantic eavesdropping attacks, semantic inference attacks, and semantic jamming attacks Once the data is written into the blockchain, it cannot be easily changed or deleted, ensuring that the semantic information is not maliciously tampered with during transmission. In addition, the decentralized nature of blockchain makes semantic information no longer rely on a single central node for storage and transmission, which reduces the risk of data being single point of failure and improves the overall security of SemCom system. Blockchain can also be used to ensure that the identities of both communication parties are authentic, thus preventing impersonation and fraud.
Model compression Model pruning, parameter quantization, low-rank decomposition, and knowledge distillation Attacks against communication bottlenecks, semantic inference attacks By employing model compression techniques, devices can effectively decrease communication costs, conserve storage space, and eliminate parameter redundancy, thereby strengthening the SemCom system to defend against against multi-party cooperation attacks.
@@
Physical-layer security Beamforming, artificial noise, relay cooperation, intelligent reflecting surface, physical-layer key generation, and physical-layer authentication Semantic eavesdropping attacks, semantic inference attacks, and semantic jamming attacks It involves utilizing the randomness of wireless channels, such as interference, fading, and noise, to ensure secure transmission.
@@ 4 Techniques to Mitigate Security and Privacy Threats in SemCom
In response to the security and privacy threats detailed in Sec. 3, this section provides defensive techniques, including data cleaning, robust learning, defensive methods of backdoor attacks, adversarial training, differential privacy, cryptography technology, block-chain technology, model compression, and physical-layer security. The detailed techniques to mitigate security and privacy threats in SemCom are shown in Tab. 4.
@@ 4.1 Data Cleaning
Dirty data can introduce noise, missing values, or inconsistencies, directly causing semantic misunderstandings and even communication failure. Fig. 5 illustrates the use of data cleaning in DL.
Figure 5: Data cleaning in DL [117], where the process ensures the quality and consistency of data through collection, cleaning, validation, training, evaluation, and serving. Data cleaning, being a fundamental step, helps improve the accuracy, robustness, and reliability of the model, enabling better overall performance in real-world applications.
The below, as shown in Table LABEL:datacleaning, the following content, based on multiple academic studies, delves into various data cleaning methods for different data types and their potential applications in SemCom.
Table 5: List of Representative Data Cleaning Schemes
[…]
@@ 4.1.2 Diverse Cleaning of Image Data
Image data cleaning refers to the process of reducing noise, correcting artifacts, and enhancing data quality in images used for SemCom. This includes techniques like noise removal through Gaussian blur and bilateral filtering, which preserve critical features while eliminating distortions. Data augmentation methods, such as rotations and scaling, further improve adaptability to diverse visual scenarios, and GANs can generate synthetic data to address training data shortages.
@@ 4.1.4 Comprehensive Optimization of Structured Data
SemCom for poisoning attack mitigation: Lotfi et al. [124] proposed VMGuard, a reputation-based framework for detecting data poisoning attacks in vehicular Metaverse systems. By integrating SemCom and subjective logic modeling, VMGuard evaluates the trustworthiness of IoT devices, ensuring reliable data collection and mitigating adversarial tampering in real-time semantic data streams.
@@ 4.2 Robust Learning
The design characteristics of FL, as shown in Fig. 6, make FL face numerous challenges in practical applications, such as noise interference, data heterogeneity, malicious clients, and low communication efficiency. These challenges not only hinder the improvement of model performance but also pose higher demands on system robustness and efficiency. The following text, some typical papers as shown in the table LABEL:robustlearning, reviews the latest advances in the robustness of federated learning, and presents possible directions for improvement as well as potential applications in SemCom.
Figure 6: Diagram of FL [125], where data is processed in a decentralized manner by multiple clients, each performing forward propagation, model updating, and backpropagation before sending gradients to the server for aggregation. This decentralized approach ensures data privacy by allowing model training on local devices without the need to share raw data. The server then aggregates the gradients from the clients to update the global model, which is sent back for further iterations, leading to a more robust and generalized model.
Table 6: List of Representative Robust Learning Schemes
[…]
@@ 4.2.3 Mitigating the Impact of Noise and Malicious Clients
Figure 7: Malicious client attacks in FL [129], where malicious clients attempt to leak sensitive information through their updates, affecting the parameter server and potentially compromising the privacy of the training process. These attacks could expose critical data such as class representatives, membership details, properties of the data, sensitive attributes, or even training inputs and labels, ultimately undermining the security and trustworthiness of the FL system.
@@ 4.2.4 Balancing Robustness and Communication Efficiency
Figure 8: Comparison of noise distributions under the expectation-based model and the worst-case model [131], where panel (a) visualizes the noise characteristics under an expectation-based model in two-dimensional space, and panel (b) depicts noise under the worst-case model in the same space. The analysis demonstrates that the expectation-based model effectively reduces the noise effects in the loss function, enhancing the global model’s robustness against communication noise. Experimental results show that the worst-case model also achieves comparable performance while maintaining strong model accuracy even in challenging communication scenarios.
Lesson 1